# Cross-Modal Reconstruction Pretraining for Ramp Flow Prediction at Highway Interchanges

### v1

@@ 5.5 Accuracy Across Ramps
To further evaluate the model performance, we conducted a micro-level analysis of the prediction accuracy for individual ramps. This section aims to examine the improvement of STDAEGWNET over its backbone network, GWNet, in capturing the dynamics of specific ramp traffic flows. We selected the results of the QiLin interchange with a 5 min sampling interval and visualized the predictions for the first day in the test set.
As shown in Figure 5, the figure intuitively presents the comparison between the predicted curves and the ground truth for STDAEGWNET and GWNet at steps 3, 6, and 12. Visually, the prediction curves of STDAEGWNET closely follow the fluctuations of the actual traffic, particularly at the peaks and troughs of traffic flow, demonstrating stronger capability in capturing traffic dynamics.
For quantitative comparison, we further calculated the performance metrics for each turning ramp over the entire test set, and the detailed results are shown in Table 8. The quantitative results clearly confirm the conclusions from the visualization analysis. From the table, it can be seen that out of 12 ramps, STDAEGWNET outperforms the original GWNet in 10 ramps in terms of overall prediction performance. Particularly for major directions with higher traffic volumes and more challenging predictions, such as “W to E” and “E to W”, STDAEGWNET shows a significant advantage, with MAE reduced by approximately 9.0% and 4.9%, respectively, and RMSE also decreased accordingly. This indicates that the model predicts complex traffic patterns more accurately and robustly.
@@
It is worth noting that for the “E to N” and “N to E” ramps, GWNet exhibits a slight advantage. Analysis of the traffic characteristics reveals that these ramps likely have relatively low traffic volumes or simpler patterns, where the deep spatiotemporal context injected by STDAE pretraining does not play a critical role, and a strong direct prediction model can already achieve good performance. Overall, the ramp-level fine-grained comparative analysis strongly demonstrates the practical value of the STDAE pretraining framework. By learning deep mapping relationships between the mainline and ramps, it significantly enhances the backbone network performance at the micro-level, especially in handling complex and high-volume traffic scenarios, with substantial improvement in predictive accuracy.
@@ 6 Conclusion
This study proposes the STDAE framework, a two-stage approach that first performs cross-modal reconstruction pretraining and then integrates the learned representations with downstream predictors to tackle the problem of ramp flow prediction in detector-scarce highway interchanges. In the pretraining stage, STDAE reconstructs ramp flows from mainline traffic sequences collected by ETC systems, enabling its spatial and temporal branches to capture heterogeneous but complementary features. After feature fusion, the framework provides downstream predictors with rich long-range contextual information.
Comprehensive experiments on multi-week ETC datasets at three real-world cloverleaf interchanges and under multiple sampling intervals confirm the effectiveness of the proposed method. We adopt GWNet as the downstream predictor, and the combined model STDAEGWNET achieves the best overall performance across MAE, MAPE, and RMSE metrics among thirteen competitive baselines on all three datasets. Moreover, compared with advanced models that rely on historical ramp flow sequences, STDAEGWNET demonstrates comparable predictive capability. Since STDAE relies solely on widely available mainline ETC data, it enables fine-grained interchange control even in the absence of ramp detectors, without requiring hardware retrofitting. Extensive ablation studies further reveal that the proposed STDAE generates informative representations that effectively capture spatial and temporal dependencies between mainline and ramp flows. The learned embeddings are architecture-agnostic and can be seamlessly injected into other sequence or graph models with minimal modifications, serving as a plug-and-play enhancement for diverse forecasting pipelines. In addition, STDAE remains robust and achieves favorable performance even under scenarios with partially missing mainline data.
The current work focuses on multi-step short-term forecasting. Extending the framework to probabilistic prediction constitutes a natural next step. External factors such as incidents and weather are not explicitly modeled, and incorporating event-aware modules could further enhance robustness.
7 Acknowledgements
The authors gratefully acknowledge the support of ITSSKY Technology Co., Ltd. for providing the data used in this study. The authors would like to express their sincere gratitude to Dr. Wenbo Lu, Ph.D. candidate Feiyang Pan, Master’s student Tao Chen, Master’s student Qianyi Hu from Southeast University, and Ph.D. candidate Yibo Chen from East China University of Science and Technology for their invaluable support and thoughtful guidance. This research work is supported by the Big Data Computing Center of Southeast University.

### v2

@@ 5.5 Accuracy Across Ramps
To further evaluate model performance, a micro-level analysis of prediction accuracy for individual ramps is conducted. This section examines the improvement of STDAEGWNET over its backbone network, GWNet, in capturing the dynamics of specific ramp traffic flows. The results from the QiLin interchange with a 5 min sampling interval are selected, and the predictions for the first day in the test set are visualized.
As shown in Figure 7, the figure intuitively presents the comparison between the predicted curves and the ground truth for STDAEGWNET and GWNet at step = 3. Visually, the prediction curves of STDAEGWNET closely follow the fluctuations of the actual traffic, particularly at the peaks and troughs of traffic flow. Four time segments from Figure 7 (a3), (b2), (c3), and (d1) are selected to illustrate the prediction results, as shown in Figure 7 (a4)-(d4). These correspond to traffic flow trends of increase, fluctuation, decrease, and uctuating increase, respectively. The blue curve represents the background ramp traffic flow, the green curve represents the GWnet prediction, and the red curve represents the prediction with STDAE added. Intuitively, adding STDAE more accurately predicts traffic flow trends of increase and decrease. It also better matches the actual traffic flow compared to GWnet in the case of traffic fluctuations.
For quantitative comparison, the performance metrics for each turning ramp over the entire test set are calculated, and the detailed results are presented in Table 8. The quantitative results clearly confirm the conclusions from the visualization analysis. From the table, it can be seen that out of 12 ramps, STDAEGWNET outperforms the original GWNet in 10 ramps in terms of overall prediction performance. Particularly for major directions with higher traffic volumes and more challenging predictions, such as “W to E” and “E to W”, STDAEGWNET shows a significant advantage, with MAE reduced by approximately 9.0% and 4.9%, respectively, and RMSE also decreased accordingly. This indicates that the model predicts complex traffic patterns more accurately and robustly.
@@
As shown in the first and last rows of Table 8, GWNet exhibits a slight performance advantage for the “E to N” and “N to E” ramps, with MAE of 2.84 and 5.74, MAPE of 50.42% and 60.29%, and RMSE of 3.78 and 7.59, respectively. Analysis of the traffic characteristics reveals that these ramps likely have relatively low traffic volumes or simpler patterns, where the deep spatiotemporal context injected by STDAE pre-training does not play a critical role, and a strong direct prediction model can already achieve good performance. Overall, the fine grained comparative analysis conducted at the ramp level clearly demonstrates the practical value of the STDAE pre-training framework. By learning deep mapping relationships between the mainline and ramps, the framework significantly enhances the performance of the backbone network at the micro level, particularly in complex traffic scenarios, leading to substantial improvements in predictive accuracy.
@@ 6 Conclusion
This study proposes the STDAE framework, a two-stage approach for ramp traffic flow prediction at highway interchanges. It first performs proxy reconstruction pre-training and then integrates the learned representations with downstream predictors. In the pre-training stage, STDAE reconstructs ramp flows from mainline traffic sequences collected by ETC systems, enabling its spatial and temporal branches to capture heterogeneous but complementary features. After feature fusion, the framework provides downstream predictors with rich long-range contextual information.
Comprehensive experiments on multi-week ETC datasets at three real-world cloverleaf interchanges and under multiple sampling intervals confirm the effectiveness of the proposed method. GWNet is adopted as the downstream predictor, and the combined model STDAEGWNET achieves the best overall performance across MAE, MAPE, and RMSE metrics among thirteen competitive baselines on all three datasets, with average ranks of 1.00, 2.11, and 1.44 on the QiLin, DanYangXinQu, and XueBu datasets, respectively. Moreover, at a 3 min sampling interval, the model achieves the lowest mean absolute errors of 4.89, 5.61, and 4.58 on the QiLin, DanYangXinQu, and XueBu datasets, respectively, demonstrating a significant accuracy advantage over other methods. Compared with advanced models that rely on historical ramp flow sequences, STDAEGWNET demonstrates comparable predictive capability while relying solely on mainline data. Since STDAE relies solely on widely available mainline ETC data, it enables fine-grained interchange control even in the absence of ramp detectors, without requiring hardware retrofitting. Extensive ablation studies further reveal that the proposed STDAE generates informative representations that effectively capture spatial and temporal dependencies between mainline and ramp flows. The learned embeddings are architecture-agnostic and can be seamlessly injected into other sequence or graph models with minimal modifications, serving as a plug-and-play enhancement for diverse forecasting pipelines. In addition, STDAEGWNET remains robust and achieves favorable performance even under scenarios with partially missing mainline data, maintaining stable competitiveness and strong generalizability across all three hub datasets.
In summary, the experimental findings validate the effectiveness of the STDAEGWNET framework. It provides a practical solution to the challenge of unavailable ramp data under real-time blind spot conditions. However, this study does not consider external influencing factors such as weather conditions, traffic incidents, or special events. The model is primarily evaluated on a limited set of highway interchanges, and its generalization to urban networks remains to be explored. Additionally, the computational cost of STDAEGWNET may pose challenges for real-time deployment at networks, and the interpretability of the learned representations for traffic management decisions could be further enhanced. Future work could address these limitations by incorporating multi-source data, extending evaluations to diverse network scenarios, optimizing model efficiency, and exploring generative ramp flow prediction under more complex and realistic conditions.